An analytical chemist is titrating 120.0mL of a 0.5200 M solution of diethylamine ((C2H5)2NH) with a 0.07300 M solution of HNO3. The pKb of diethylamine is 2.89. Calculate the pH of the base solution after the chemist has added 395.6mL of HNO3 solution to it.
Answer #1

Ansues:- SzHEANH] = 0.5200M, 120.0m FHNO3) = 0.07300M 398-6 oko 2.89 -pka = 14-pko: 14-2-89 _ =11.11 (SzHs), MH + HO-945), NyAnswer:-

This question is solved by using the simple concept of calculation of pH of buffer solution using Henderson Hasselbach equation.

The answer is given in the image,